OpenStudy (anonymous):

A line passes through the point (–2, 4), and its y-intercept is (0, –6). Can someone put this into slope intercept form for me? i am not sure how to work this out. i need it to find a line that is perpendicular to it for on another question.

OpenStudy (whpalmer4):

Find the slope of the line using the formula \[m = \frac{y_2 - y_1}{x_2-x_1}\]You have the two points needed. Once you've found the value of \(m\), you can write the slope-intercept form by plugging it into \[y=mx+b\]where \(b\) is the \(y\) value of the y-intercept.

OpenStudy (whpalmer4):

to clarify my first post: you have two known points \((x_1,y_1) = (-2,4)\) and \((x_2,y_2) = (0,-6)\) (it doesn't matter which point you choose as \((x_1,y_1)\), just that you be consistent in the problem)

OpenStudy (anonymous):

yeah i meant i need it in the y=mx+b kinda form

OpenStudy (whpalmer4):

and I've given the steps you need to follow to accomplish that

OpenStudy (whpalmer4):

or a set of steps to follow to accomplish that. not the only possible route...

OpenStudy (whpalmer4):

the slope of your perpendicular line will be \(-1/m\) where \(m\) is the slope of this line

OpenStudy (anonymous):

so the slope of this line is positive 1?

OpenStudy (whpalmer4):

mmm...don't think so.  Does that look like a line that goes up 1 unit every 1 unit you move to the right?

OpenStudy (anonymous):

rise=5 and the run=1 so the slope is 5/1 which means my perpendicular lines slope will be -1/5 and if my points for my other line are (5,-4) than the equation would be y=1/5x-5 is this correct.

OpenStudy (whpalmer4):

No, the rise is not 5. A positive value of the rise means going to the right causes the line to go up, not down.

OpenStudy (whpalmer4):

\[m = \frac{-6-4}{0-(-2)} = \frac{-10}{2} = -5\] or if you chose the points in the other order: \[m = \frac{4-(-6)}{-2-0} =\frac{10}{-2} = -5\]

OpenStudy (whpalmer4):

Any time you have to find the slope, you ought to make a quick sketch of the points and make sure your answer is reasonable. If you have any understand of the notion of slope, it's hard to look at my diagram and conclude that either 1 or 5 are reasonable values for the slope of that line. This is a very important concept, so now's the time to make sure your understanding is correct!